Name: 1. Prove that if N is a subgroup of index 2 in a group G, then N is normal in G 2. Let N < SI consists of all those permutations ơ such that o(4)-4. Is N nonnal in sa? 3. Let G be a finite group and H a subgroup of G of order . If H is the only subgroup of G of order n, then is normal in G 4. Let G be a group, and kt C-ra e Glar-ra for all reG} be the center of G. Show that C is normal subgroup of G 5. Consider the subgroups < 6 > and < 30 > of Z·Show that < 6 >/< 30 > 6. Consider the group G = Z16 with 11 = < 9 > and K =< 18 >. Zs. (a) List the cosets in G/H, shoning the elements in each coset (b) List the cosets in G/K, showing the elements in each coset (e) List the cosets in H/K, showing the elements in each coset (d) List the cosets in (G/K)/(H/K), showing the elements in each coset e Show that (G/H) (G/K)/(H/K). [Define an isomorphism between G/H and (G/k)(H/K) 7. Let ф : G-, G , where G is any group, be given by Og)-9-1, for g e G. Determine whether the map φ is a hornomorphisan. 8. Let ф : G- G, be a honomorphisin with kernel 11 and let a E G. Iet S-{r e G e(1)- H@ and HaS s). φ(a)]. Prove S = Ha. (Hint Show that S 9. List all elements of Zs x Z6 and the elements of the cyelie group < (0,1) >. Then compute the factor group(Z x 26)/< (0,1) > 10. Let φ: Zis·→ Z12 be the limomorphismi wiereo(1)-10. (a) Find K-ker() (b) List the cosets in Zjs/K showing the elements in each coset (c) Find ф1z,a]. 0 495
Answer Question 5 .
Answer #1

Let G=<6> and H=<30> . Since G is an abelian group, H is a normal subgroup of G . Therefore G/H is a group ( with elements of the form g+H , where g \in G ).

The elements of G/H are:

0+H=H=\{0,\pm 30,\pm 60,\pm 90,...\},

6+H=\{...,-84,-54,-24,6,36,66,96,...\},

12+H=\{...,-78,-48,-18,12,42,72,102,...\},

18+H=\{...,-72,-42,-12,18,48,78,108,...\} and

4H-...,-66,-36, -6, 24,54, 84, 114,....

See that H is the zero element in G/H .

We form the mapping \phi : \mathbb{Z}_5\rightarrow G/H by \phi(\overline{k})=6k+H for 0\leqslant k\leqslant 4, k \in \mathbb{Z} .

So \overline{0} \in \mathbb{Z}_5 is mapped to the zero element in G/H and \phi (m+n)=6(m+n)+H=(6m+H)+(6n+H)=\phi(m)+\phi(n) from the rules of coset addition.   So \phi is a homomorphism.

The map \psi:G/H\rightarrow \mathbb{Z}_5 given by \psi(6k+H)=\overline{k} for 0\leqslant k\leqslant 4, k \in \mathbb{Z} is the inverse of \phi and this shows that \phi is bijective.

Hence \phi is an isomorphism and so <6>/<30>\cong \mathbb{Z}_5 .

P.S.: Please upvote if you have found this answer helpful.